OpenStudy (anonymous):

What is the solution to |x-10|=3x-2?

OpenStudy (anonymous):

But don't I have to make them two separate equations or something, that's what confuses me?

OpenStudy (anonymous):

if \(x>10\) then \[|x-10|=x-10\] solve \[x-10=3x-2\]

OpenStudy (anonymous):

you get \[-10=2x-2\\ -12=2x\\ x=-6\] but this is NOT a solution because we were assuming that \(x>10\) and \(-6\) is less than \(10\)

OpenStudy (anonymous):

oh i made a mistake doh

OpenStudy (anonymous):

\[x-10=3x-2\\ -10=2x-2\\ -8=2x\\ x=-4\] still NOT A SOLUTION

OpenStudy (anonymous):

then if \(x<10\) you have \(|x-10|=10-x\) solve \[10-x=3x-2\]

OpenStudy (anonymous):

\[10-x=3x-2\\ 10=4x-2\\ 12=4x\\ x=3\]

OpenStudy (anonymous):

I'm really lost sorry I just don't understand what you're doing?

OpenStudy (anonymous):

and that IS a solution since \(3<10\)

OpenStudy (anonymous):

lets go slow

OpenStudy (anonymous):

you have to get rid of the absolute values when you solve this equation

OpenStudy (anonymous):

in order to do that, you have to know whether \[x-10\] is positive or negative if \(x-10\) is positive, then \(|x-10|=x-10\) is that much clear?

OpenStudy (anonymous):

Wait so what are the two equations?

OpenStudy (anonymous):

we are not there yet the two equations to solve are \[x-10=3x-2\] or \[-(x-10)=3x-2\] but we have to know something else first

OpenStudy (anonymous):

(after consulting my algebra book) | x-10 | = 3x-2 x+10>3x-2 -x -x 10>2x-2 +2 +2 12/2>2x/2 6>x

OpenStudy (anonymous):

@BewaretheDragon this is not an inequality

OpenStudy (anonymous):

and that answer is wrong in any case

OpenStudy (anonymous):

Okay so what else do I need to know?